A d.c series motor is connected across a 460 - V supply runs at 500 r/min when the current is 40 A . The total resistance of the armature and field circuits is 0.6 ohms. calculate the torque on the armature in Newton meters.
1
Expert's answer
2021-03-11T06:56:21-0500

"Input \space voltage=460V"


"Speed(N) =500rpm\implies \newline\omega=2\pi N/60=52.36 \space rad/sec"


"Current(Ia) =40A"


"Total \space resistance\space of armature \space and \space field=\newline Rse+Ra=0.6\Omega"

"We \space have, 
V=Eb+Ia(Ra+Rse) \newline
Hence , Back \space emf (Eb) =460-40\times0.6=436V"

"We \space have, \newline 
EbIa=T\omega\implies\newline Torque(T) \newline=EbIa/\omega=436\times40/52.36= 333.08Nm"